OpenStudy (anonymous):

SOLVE THE EQUATION FOR X. ln e^x = -3

OpenStudy (anonymous):

lne^x=-3 literally means x=-3

OpenStudy (anonymous):

lne^x = x

OpenStudy (anonymous):

Therefore, x = -3

OpenStudy (anonymous):

\[ \ln e^x=-3\\ e^{\ln e^x}=e^{-3}\\ e^x=e^{-3}\\ x=-3 \]

OpenStudy (anonymous):

lne^x read as a sentence: "what do I need to raise e to in order to get e^x"